OpenStudy (anonymous):

QUICK EASY YES OR NO QUESTION PLEASE HELP. Does sine 7pi/3 equal root3/2?

OpenStudy (across):

yes

OpenStudy (anonymous):

\[\sin(7pi/3)=\sin(pi/3)=\sqrt{3}/2\]